# Burgers equation from non-thermal stationary states in nearly-integrable gases

### v1

@@ II.2 GHD-Boltzmann equation
The dynamics of integrable systems in the presence of integrability-breaking effects is given by GHD-Boltzmann equation
@@
where $\mathcal{I}[\rho_{\mathrm{p}}]$ is the collision operator responsible for interaction between the observed system $\hat{H}_{\rm int}$ and bath $\hat{H}^{(b)}_{\rm int}$. We assume that the bath system is much larger and in consequence, does not evolve in time. Equation (15) describes the system only and thus does not conserve momentum and energy. However, it does conserve particle number which is reflected in collision integral
@@ III Athermal stationary states
Athermal states relevant to the present context were first found in the setup of coupled Lieb-Liniger gases [58]. In this work, thanks to the recent development of gBBGKY [30], we present arguments which hold for arbitrary weakly coupled integrable systems. The gBBGKY assumes that the system and bath interact through long-range density-density interaction given by
@@
where $\xi$ is the characteristic range of the potential, $V_{0}$ measures it’s strength and function $\varphi$ gives the interaction profile. We will assume here that the bath does not evolve, which can be realized for instance by a large density imbalance between the two systems. The aforementioned kinetic framework predicts that the dynamics of the system is given by Boltzmann-like kinetic equation.
$\partial_{t}\rho_{\mathrm{p}}=\mathcal{I}_{2p}+\mathcal{O}(\xi^{-1}),$ (18)
where collision integral $\mathcal{I}_{2p}$ describes two-particle scattering processes, which will be discussed in the following. There are important corrections, written above as $\mathcal{O}(\xi^{-1})$, which involve collision operators responsible for three and more particle collisions. These processes are relevant on timescales longer than $\mathcal{I}_{2p}$ and eventually thermalize the system. In what follows, we neglect them focusing on timescales relevant for $\mathcal{I}_{2p}$. Let us also note that the same picture emerges from Fermi’s Golden Rule treatment of integrability breaking as developed in [36, 20]. There, the collsion operators involve terms with more and more particle collision processes, with the two-particle scatterings dominating in the system-bath setup.
@@
$\mathcal{I}[\rho_{\mathrm{p}}](\lambda)]$ $=\int{\rm d}\lambda^{\prime}{\rm d}\lambda_{b}{\rm d}\lambda_{b}^{\prime}\mathcal{W}(\lambda,\lambda_{b};\lambda^{\prime},\lambda_{b}^{\prime})\delta(k)\delta(\omega)\times$ (19)
$\left(\rho_{\mathrm{p}}(\lambda)\rho_{\mathrm{p}}^{(b)}(\lambda^{\prime})-\rho(\lambda^{\prime})\rho_{\mathrm{p}}^{(b)}(\lambda)\right),$
where conservation laws are imposed by Dirac deltas with $k=k(\lambda)+k_{b}(\lambda_{b})-k(\lambda^{\prime})-k_{b}(\lambda^{\prime}_{b})$ and $\omega=\omega(\lambda)+\omega_{b}(\lambda_{b})-\omega(\lambda^{\prime})-\omega_{b}(\lambda^{\prime}_{b})$, where $k,k_{b}$ are momenta of system and bath particles, the same goes for energies $\omega,\omega_{b}$. The quantity $\mathcal{W}(\lambda,\lambda_{b};\lambda^{\prime},\lambda_{b}^{\prime})$ is naturally related to a rate of scattering $\lambda^{\prime}+\lambda^{\prime}_{b}\to\lambda+\lambda_{b}$.
Importantly, we assume that dispersion relations in both systems are not necessarily the same (meaning $\omega_{b}\neq\omega$ and $k_{b}\neq k$), which in the context of classical particles can be realized for instance by different masses, or different length of the rods (in the case of hard rod gas). Non-equal dispersions generically prohibit the existence of athermal state and thermal states are the only fixed points of (19).
The situation of equal dispersion happens when both gases consist of identical free classical particles $k(\lambda)=\lambda$ and $\omega(\lambda)=\lambda^{2}/2$, the momentum-energy conservation laws imply that scattering amounts to permutation of velocities. We readily find the stationary state which is
@@
Obviously, if $\rho^{(b)}$ is athermal then the state above is athermal as well. This is the simplest example of dynamics leading to non-thermal stationary state.
Remarkably, this changes in the grazing limit of the equation (19), which amounts to restricting possible processes to only these, where the momentum exchange is small. In that limit, the collision integral acquires the Landau form
$\mathcal{I}^{\rm gr}[\rho_{\mathrm{p}}](\lambda)$ $=-\partial_{\lambda}\int{\rm d}\lambda_{b}\tilde{W}(\lambda,\lambda_{b})\delta(v(\lambda)-v_{b}(\lambda_{b}))\times$ (21)
$\rho_{\mathrm{p}}(\lambda)\rho_{\mathrm{p}}^{(b)}(\lambda_{b})\left[\epsilon^{\prime}(\lambda)-\frac{k^{\prime}(\lambda)}{k^{\prime}_{b}(\lambda_{b})}(\epsilon^{(b)})^{\prime}(\lambda_{b})\right],$
with $\tilde{\mathcal{W}}(\lambda,\lambda_{b})$ directly related to collision rate $\mathcal{W}$. The stationary state fulfils two sets of equations
@@
This stationary state appears even though two-body scattering does not lead to simple momentum exchange. Let us stress that the solution to the equation above is athermal in the case when bath is not in a thermal state. In [58] such athermal states were found in the setup, where both the system and bath are described by the Lieb-Liniger models.
In the generic interacting integrable models, the dispersion relation is a dynamic quantity, it depends on the state of the system which in turn reflects e.g. the interactions and the density. Therefore, two such systems that are weakly coupled naturally provide a physical realisation in which two body scatterings in (1+1)d lead to momentum redistribution. In this case we expect the stationary state described by (22).
@@
Table 2: Conditions for possible existence of athermal states in two-particle scattering processes described by collision integral (19). Athermal states are excluded only if the dispersions of the systems $H_{\rm int}$, $H_{\rm int}^{(b)}$ are different, and the grazing limit of small momentum exchange is absent.
In what follows we present different treatments of weakly coupled integrable systems. We start with the most concrete case of coupled Lieb-Linger models, which can be studied using Fermi’s Golden Rule. This analysis is later generalized to arbitrary weakly coupled integrable models using the gBBGKY hierarchy. Lastly, we propose a simplified model of stochastic two-body collisions which offers particularly simple structure of collision integral while capturing all the relevant features of the dynamics. It also has the advantage of efficient numerical simulation and exact expressions for diffusion constants.
@@ III.1 Fermi’s Golden Rule
We start with the Fermi’s Golden rule treatment of nearly integrable quantum models. Within this approximation, the collision integral involves matrix elements of the perturbing operator. In our case this origins from the integrability breaking term in the Hamiltonian which we assume couples to the local particles’ density operator $\hat{q}_{0}(x)$. However, we can be slightly more general and assume that it couples to density $\hat{q}_{j}(x)$ of any of the conserved charges $\hat{Q}_{j}=\int{\rm d}x\,\hat{q}_{j}(x)$.
The collision integral takes then the following form
@@
where the back-flow operator $\mathbb{F}$ describes the response of the whole state on the scattering process described by the bare collision integral $Q_{0}[\rho_{\rm p}]$. The latter is given by the usual form from the Master equation,
@@
with the transition rate $w(\lambda,\mu)$. The transition rate is computed from the Fermi’s golden rule and involves matrix elements of the perturbing operator $\hat{q}_{j}(x)$. Typically, the transition rate obeys a detailed balance relation,
@@
To illustrate a situation in which athermal stationary states exist we consider a two coupled 1d systems, see Fig. 1 with total Hamiltonian
$H=H_{1}+H_{2}+V_{\rm pert}.$ (26)
We assume that $H_{i}$ are integrable, for example are given by the Lieb-Liniger model, and $V_{12}$ breaks the integrability. In this formulation it is natural to treat system and bath on equal footing which is reflected in a slightly different notation than in eq. (4). Experimentally relevant choice of $V_{12}$ is the long-range interaction potential between two tubes that couples to local densities
$V_{\rm pert}=\int{\rm d}x_{1}{\rm d}x_{2}\,V(x_{1}-x_{2})\hat{q}_{i}(x_{1})\hat{q}_{j}(x_{2}).$ (27)
@@
In the homogeneous system, the quasi-particles density in the two tubes evolves according to the set of equations
@@
$w_{12}(\lambda,\mu)$ $=\hat{V}^{2}(k)|F_{i}^{(1)}(\lambda,\mu))|^{2}S_{j}^{(2)}(-k,-\omega),$ (29)
$w_{21}(\lambda,\mu)$ $=\hat{V}^{2}(k)|F_{j}^{(2)}(\lambda,\mu))|^{2}S_{i}^{(1)}(-k,-\omega).$ (30)
Here $F_{i}^{(1)}(\lambda,\mu)$ is the one particle-hole form factor of the operator $\hat{q}_{i}$ evaluated in the first tube. The rapidity of the particle and the hole are $\lambda$ and $\mu$ respectively. With $k\equiv k(\lambda)-k(\mu)$ and $\omega\equiv\omega(\lambda)-\omega(\mu)$ we denote momentum and energy of the particle-hole pair. $\hat{V}(k)$ is the Fourier transform of the interaction potential from eq. (27). Finally, $S_{j}^{(2)}(k,\omega)$ is the dynamic structure factor in the second tube, namely Fourier transform of the connected two point function $\langle\hat{q}_{j}^{(2)}(x,t)\hat{q}_{j}^{(2)}(0)\rangle^{c}$ of density $\hat{q}_{j}$. In principle, there are also contributions to the transition rates that involve form-factors with larger number of particle-hole pairs. As shown in [52] their contribution is subleading in the momentum transferred between the tubes and is important at larger time scales. Importantly, the higher particle-hole processes are responsible for ultimate thermalization of the system [20].
@@
$w(\lambda,\mu)=\hat{V}^{2}(k)|F_{i}^{(1)}(\lambda,\mu))|^{2}S_{j}^{(2)}(-k,-\omega),$ (31)
@@
Note also, that for illustration purposes the derivation of the collision integral was presented in the case of two coupled integrable models with density imbalance. However, the only information that we need of the bath is its dynamic response function $S_{j}^{(2)}$. Instead, for the description of the system, its integrability is crucial as it provides the quasi-particle picture.
@@
Following [52] let us discuss some specific cases. First, if we specialize to a thermal state, then $\epsilon_{2}^{\prime}(\lambda)=\beta\omega_{2}^{\prime}(\lambda)$ and $\mathcal{F}_{2}(v)=\beta v$. The stationary state of the system is then thermal with the same inverse temperature $\beta$. The second special case is when states and Hamiltonians in both tubes are the same. In this cases the collision integral (at the 1 particle-hole level) is exactly zero. However, such situation cannot occur in the self-diffusion problem since the densities, and therefore states, are different by construction. Finally, in the Tonks-Girardau limit of the Lieb-Liniger model the dressing is absent. Then the effective velocities are state independent and equal to $h_{1}$. The stationarity condition simplifies to $\epsilon_{1,0}^{\prime}(\lambda)=\epsilon_{2,0}^{\prime}(\lambda)$. Crucially, in the Tonks-Girardeau gas the form-factors with larger number of particle-hole excitations vanish and the athermal states are stationary beyond the small momentum approximation.
@@ III.2 Generalized BBGKY hierarchy
In this section we take a more general perspective on the integrability breaking and employ recently developed generalization of BBGKY hierarchy for the dynamics of correlation functions [30]. We specialize to the system described by the Hamiltonian (26) and assume homogenous states. Importantly, the interaction potential is assumed to be long-ranged and given by the form (17). The full hierarchy is exposed in [30] and involves correlation functions of all orders. To understand the appearance of athermal states it is sufficient however to truncate the hierarchy at the level of two-particle (system-bath) correlations $g_{\lambda_{1},\lambda_{2}}{(x_{1},x_{2})}$. Such correlations capture two-body scattering physics discussed earlier. This means that we have to deal with three dynamical equations (we still use the notation where index $1$ refers to the system and $2$ refers to the bath)
@@
Here the flux Jacobian $\mathcal{A}$ and the susceptibility matrix $\mathcal{C}$ of eqs. (12) and (13) are evaluated in the quasi-particle basis, see Appendix A for the details.
The hierarchy can be integrated in time, leading to coupled equations of the form of (28). In Appendix A we go through this calculation, which follows standard procedure of deriving collision operators from the BBGKY hierarchy. The resulting collision integrals take the Landau form in the grazing limit. The explicit form is given in (116). The resulting collision operator, upon specifying to Lieb-Linger model exactly agrees with Fermi’s Golden Rule result. The athermal states featured by it are exactly of the form (22).

### v2

@@ II.2 GHD-Boltzmann equation
We will assume here that the dynamics of integrable systems in the presence of integrability-breaking effects is given by GHD-Boltzmann equation [36, 38]
@@
where $\mathcal{I}[\rho_{\mathrm{p}}]$ is the collision operator responsible for interaction between the observed system and the bath. We indicated the dependence on state in effective velocity and diffusion operator by putting $\rho$ in the lower index. Moreover, we assume that the bath system is much larger and in consequence, does not evolve in time. Equation (15) describes the system only and thus does not conserve momentum and energy. However, it does conserve particle number, which is reflected in collision integral
@@ III Athermal stationary states
Athermal states relevant to the present context were first found in the setup of coupled Lieb-Liniger (LL) gases [68]. In this work, thanks to the recent development of gBBGKY [48], we present arguments which hold for arbitrary weakly coupled integrable systems. In gBBGKY for coupled integrable systems we assume that the system and bath interact through long-range density-density interaction given by
@@
where $\xi$ is the characteristic range of the potential, $V_{0}$ measures its strength and function $\varphi$ gives the interaction profile. Importantly, the parameter $\xi$ should be thought of as a large parameter in the theory in order to provide a mechanism for truncation of the gBBGKY hierarchy. Moreover, we will assume here that the bath does not evolve, which can be realized for instance by a large density imbalance between the two systems. The aforementioned kinetic framework predicts that the dynamics at sufficiently late times of the system is given by Boltzmann-like kinetic equation.
$\partial_{t}\rho_{\mathrm{p}}(\lambda)=\mathcal{I}_{2p}(\lambda)\left(1+\mathcal{O}(\xi^{-1})\right),$ (18)
where collision integral $\mathcal{I}_{2p}$ describes two-particle scattering processes, which will be discussed in the following. There are important corrections, written above as $\mathcal{O}(\xi^{-1})$, which involve collision operators responsible for three and more particle collisions. These processes are relevant on timescales longer than $\mathcal{I}_{2p}$ and eventually thermalize the system. In what follows, we neglect them focusing on timescales relevant for $\mathcal{I}_{2p}$. Let us also note that the same picture emerges from Fermi’s Golden Rule (FGR) treatment of integrability breaking as developed in [36, 38]. There, the collision operator involve terms with more and more particle collision processes, with the two-particle scatterings dominating in the system-bath setup. The exact equivalence of between FGR and gBBGKY was established in [48].
@@
$\mathcal{I}[\rho_{\mathrm{p}}](\lambda)$ $=\int{\rm d}\lambda^{\prime}{\rm d}\lambda_{b}{\rm d}\lambda_{b}^{\prime}\mathcal{W}(\lambda,\lambda_{b};\lambda^{\prime},\lambda_{b}^{\prime})\times$ (19)
$\delta(k)\delta(\omega)\left(\rho_{\mathrm{p}}(\lambda)\rho_{\mathrm{p}}^{(b)}(\lambda^{\prime})-\rho_{\mathrm{p}}(\lambda^{\prime})\rho_{\mathrm{p}}^{(b)}(\lambda)\right),$
where conservation laws for a scattering process are imposed by Dirac deltas with $k=k(\lambda)+k_{b}(\lambda_{b})-k(\lambda^{\prime})-k_{b}(\lambda^{\prime}_{b})$ and $\omega=\omega(\lambda)+\omega_{b}(\lambda_{b})-\omega(\lambda^{\prime})-\omega_{b}(\lambda^{\prime}_{b})$. In general, we assume that quasiparticle with quasimomentum $\lambda$ carry momentum $k(\lambda)$ and energy $\omega(\lambda)$. In our notation $k,k_{b}$ are momenta of system and bath particles, respectively, the same goes for energies $\omega,\omega_{b}$. The quantity $\mathcal{W}(\lambda,\lambda_{b};\lambda^{\prime},\lambda_{b}^{\prime})$ is naturally related to a rate of scattering $\lambda^{\prime}+\lambda^{\prime}_{b}\to\lambda+\lambda_{b}$.
Importantly, we assume that dispersion relations in both systems are not necessarily the same (meaning $\omega_{b}\neq\omega$ and $k_{b}\neq k$), which in the context of classical particles can be realized for instance by different masses, or different length of the rods (in the case of hard rod gas). Non-equal dispersions generically prohibit [68] the existence of athermal state and thermal states (given by Maxwell-Boltzmann distributions) are the only fixed points of (19).
The complementary situation of equal dispersion happens when both gases consist of identical free particles with $k(\lambda)=\lambda$ and $\omega(\lambda)=\lambda^{2}/2$. The momentum-energy conservation laws imply then that scattering amounts to permutation of velocities meaning that $\lambda=\lambda_{b}^{\prime}$ and $\lambda_{b}=\lambda^{\prime}$. We readily find the stationary state in such case which is
@@
Obviously, if $\rho_{\mathrm{p}}^{(b)}$ is athermal then the state above is athermal as well. The case of two classical gases with equal dispersions is the simplest example of dynamics leading to non-thermal stationary state.
Remarkably, the conclusion about lack of non-thermal fixed points in the case of non-equal dispersions changes in the grazing limit of the equation (19), which amounts to restricting possible scattering processes to only these, where the momentum exchange is small. In that limit, the collision integral acquires the form of Landau-like kinetic equation [49, 75]
$\mathcal{I}^{\rm gr}[\rho_{\mathrm{p}}](\lambda)$ $=-\partial_{\lambda}\int{\rm d}\lambda_{b}\tilde{\mathcal{W}}(\lambda,\lambda_{b})\delta(v(\lambda)-v_{b}(\lambda_{b}))\times$ (21)
$\rho_{\mathrm{p}}(\lambda)\rho_{\mathrm{p}}^{(b)}(\lambda_{b})\left[\epsilon^{\prime}(\lambda)-\frac{k^{\prime}(\lambda)}{k^{\prime}_{b}(\lambda_{b})}\epsilon_{b}^{\prime}(\lambda_{b})\right],$
with $\tilde{\mathcal{W}}(\lambda,\lambda_{b})$ directly related to collision rate $\mathcal{W}$. For the precise definition of grazing limit as well as expression for $\tilde{\mathcal{W}}$ see Appendix A.
The stationary state of (22) fulfils a set of two equations
@@
Let us emphasize that this stationary state appears even though two-body scattering does not lead to a simple momentum exchange. We also stress that the solution to the equation above is athermal in the case of athermal bath. In [68] such athermal states were found in the setup, where both the system and bath are described by the LL models.
In the generic interacting integrable models, the dispersion relation is a dynamic quantity, it depends on the state of the system which in turn reflects e.g. the interactions and the density. Therefore, two such systems that are weakly coupled with potential that allows only for small exchange of momentum naturally provide a physical realisation in which two body scatterings in (1+1)D lead to momentum redistribution. In this case we expect the stationary state described by (22).
Summing up, the existence of possible non-thermal states depends on dispersion relations in both systems and on grazing limit of collision operator, cf. Tab. 2. Lastly, let us mention, that the found stationary states become thermal states in the case of thermal bath.
@@
Table 2: Conditions for possible existence of athermal states in dynamics driven by two-particle scattering processes. Athermal states are excluded only if the dispersions of the systems $H_{\rm int}$, $H_{\rm int}^{(b)}$ are different, and the grazing limit of small momentum exchange is absent.
In what follows we go beyond the case of classical particles described by (19) and present different treatments of weakly coupled integrable systems, quantum or classical. We start with the most concrete case of coupled LL models, which can be studied using FGR. This analysis is later generalized to arbitrary weakly coupled integrable models using the gBBGKY hierarchy. Lastly, we propose a simplified model of stochastic two-body collisions which offers particularly simple structure of collision integral, while capturing all the relevant features of the dynamics. It also has the advantage of efficient numerical simulation and analytical expressions for diffusion constants.
@@ III.1 Fermi’s Golden Rule
We start with the FGR treatment of nearly-integrable quantum models. Within this approximation, the collision integral involves matrix elements of the perturbing operator. In our case this origins from the integrability breaking term in the Hamiltonian which we assume couples to the local particles’ density operator $q_{0}(x)$. However, we can be slightly more general and assume that it couples to density $q_{j}(x)$ of any of the conserved charges $Q_{j}=\int{\rm d}x\,q_{j}(x)$.
The collision integral for two-particle processes takes then the following form
@@
where the backflow operator (9) describes the response of the whole state on the scattering process described by the bare collision integral $\mathcal{I}_{0}[\rho_{\mathrm{p}}]$. The latter is given by the usual form of the Master equation,
@@
with the transition rate $w(\lambda,\mu)$ and density of holes defined as $\rho_{\rm h}=\rho_{\rm t}-\rho_{\rm p}$. The transition rate is computed from the FGR and involves matrix elements of the perturbing operator $q_{j}(x)$. Typically, the transition rate obeys a detailed balance relation,
@@
To illustrate a situation in which athermal stationary states exist we consider a two coupled 1D systems, see Fig. 1(b), with total Hamiltonian
$H=H_{1}+H_{2}+V_{12}.$ (26)
We assume that $H_{i}$ are integrable, for example are given by the LL model [76, 77], and $V_{12}$ breaks the integrability. In calculations, it is natural to treat system and bath on equal footing which is reflected in a slightly different notation than in Eq. (4). Experimentally relevant choice of $V_{12}$ is the long-range interaction potential between two tubes that couples to local densities
$V_{12}=\int{\rm d}x_{1}{\rm d}x_{2}\,V(x_{1}-x_{2})q_{i}(x_{1})q_{j}(x_{2}).$ (27)
@@
In the homogeneous system, the quasiparticles density in the two tubes evolves according to the set of equations
@@
$w_{12}(\lambda,\mu)$ $=\tilde{V}^{2}(k)|F_{i}^{(1)}(\lambda,\mu))|^{2}S_{j}^{(2)}(-k,-\omega),$ (29)
$w_{21}(\lambda,\mu)$ $=\tilde{V}^{2}(k)|F_{j}^{(2)}(\lambda,\mu))|^{2}S_{i}^{(1)}(-k,-\omega).$ (30)
Here $F_{i}^{(1)}(\lambda,\mu)$ is the one particle-hole form factor of the operator $q_{i}$ evaluated in the first tube. The rapidity of the particle and the hole are $\lambda$ and $\mu$ respectively. With $k\equiv k(\lambda)-k(\mu)$ and $\omega\equiv\omega(\lambda)-\omega(\mu)$ we denote momentum and energy of the particle-hole pair. $\tilde{V}(k)$ is the Fourier transform of the interaction potential from Eq. (27). Finally, $S_{j}^{(2)}(k,\omega)$ is the dynamic structure factor in the second tube, namely Fourier transform of the connected two point function $\langle q_{j}^{(2)}(x,t)q_{j}^{(2)}(0)\rangle^{c}$ of density $q_{j}$. In principle, there are also contributions to the transition rates that involve form-factors with larger number of particle-hole pairs. As shown in [68] their contribution is subleading in the momentum transferred between the tubes and is important at larger time scales. Importantly, the higher particle-hole processes are responsible for ultimate thermalization of the system [38].
@@
$w(\lambda,\mu)=\tilde{V}^{2}(k)|F_{i}^{(1)}(\lambda,\mu))|^{2}S_{j}^{(2)}(-k,-\omega),$ (31)
@@
Note also, that for illustration purposes the derivation of the collision integral was presented in the case of two coupled integrable models with density imbalance. However, the only information that we need of the bath is its dynamic response function $S_{j}^{(2)}$. Instead, for the description of the system, its integrability is crucial as it provides the quasiparticle picture.
@@
Following [68] let us discuss some specific cases. First, if we specialize to a thermal state, then $\epsilon_{2}^{\prime}(\lambda)=\beta\omega_{2}^{\prime}(\lambda)$ and $\mathcal{F}_{2}(v)=\beta v$. The stationary state of the system is then thermal with the same inverse temperature $\beta$. The second special case is when states and Hamiltonians in both tubes are the same. In this cases the collision integral (at the 1 particle-hole level) is exactly zero. However, such situation cannot occur in the self-diffusion problem since the densities, and therefore states, are different by construction. Finally, in the Tonks-Girardau limit of the LL model the dressing is absent. Then the effective velocities are state independent and equal to $h_{1}$. The stationarity condition simplifies to $\epsilon_{1,0}^{\prime}(\lambda)=\epsilon_{2,0}^{\prime}(\lambda)$. Crucially, in the Tonks-Girardeau gas the form-factors with larger number of particle-hole excitations vanish and the athermal states are stationary beyond the small momentum approximation, due to equal dispersions in both systems.
@@ III.2 Generalized BBGKY hierarchy
In this section we take a more general perspective on the integrability breaking and employ recently developed gBBGKY hierarchy for the dynamics of correlation functions [48]. We specialize to the system described by the Hamiltonian (26) and assume homogenous states. Importantly, the interaction potential is assumed to be long-ranged and given by the form (17). The full hierarchy is exposed in [48] and involves correlation functions of all orders. To understand the appearance of athermal states it is sufficient however to truncate the hierarchy at the level of two-particle (system-bath) connected correlation function $g_{\lambda_{1},\lambda_{2}}{(x_{1},x_{2})}$. Such correlation is sufficient to capture two-body scattering physics discussed earlier. Obviously, there are more terms and equations in the hierarchy and they describe effects ultimately leading to thermalization. However, this will happen on a longer timescale and we neglect these effects in present discussion. On the level of hierarchy, this means that we have to deal with three dynamical equations (for the purpose of this section, we use the notation where index $1$ refers to the system and $2$ refers to the bath and quasimomentum dependence is indicated in lower index)
@@
The hierarchy can be integrated in time, as in standard derivations of collision kernels from BBGKY hierarchy [48, 50], leading to coupled equations of the form of (28). In Appendix B we go through this calculation. The resulting collision integrals take the Landau form in the grazing limit. The explicit form is given in (122). The resulting collision operator, upon specifying to LL model exactly agrees with FGR result. The athermal states featured by it are exactly of the form (22). The approach of gBBGKY is the most versatile as it captures both classical and quantum systems. In particular, it allowed for derivation of collision term for long-range interacting hard rods, successfully tested against exact molecular dynamics in [48]. In general, when specified to classical statistics the resulting collision integral exactly takes the form of (22) with
$\tilde{\mathcal{W}}(\lambda,\lambda_{b})=2\pi^{2}\int{\rm d}k|k|\tilde{V}^{2}(k)(k_{b}^{\prime}(\lambda_{b}))^{2}$ (37)
where $\tilde{V}(k)$ denotes Fourier transform of interaction potential (17). For more details see Appendix B.